Circular No. 7613

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET P/2001 H5 (NEAT) E. F. Helin, S. Pravdo, and K. Lawrence, Jet Propulsion Laboratory,

report the discovery of a comet (discovery observation given below) on CCD images taken with the 1.2-m Oschin Schmidt telescope at Palomar in the course of the NEAT program. CCD images taken on Apr. 25.0 UT by M. Tichý and M. Kočer at Kleť (0.57-m f/5.2 reflector) show the diffuse object to have a coma diameter of 9′′. C. E. Lopez reports that CCD observations obtained on Apr. 25.2 at El Leoncito (0.5-m f/7.5 double astrograph) also show the object to be diffuse. Additional astrometry (including LINEAR prediscovery observations on Mar. 20 identified by B. G. Marsden) and orbital elements (T = 2001 Jan. 27.0 TT, q = 2.390 AU, i = 8^{o}.4, P = 15.0 yr) are given on MPEC 2001-H37.

2001 UT  α_{2000}  δ_{2000}  m_{1}
Apr. 24.35457  14^{h}43^{m}27.^{s}88  −29^{o}40′59′′.0  16.8

SUNGRAZING COMETS Further to IAUC 7612, D. Hammer and D. Biesecker report measure-

ments by Hammer for several additional sungrazing comets found with SOHO, with all but C/2001 E1 being of Kreutz type. The reduced observations and orbital elements by B. G. Marsden appear on the MPECs cited below. Comets C/2001 F2, C/2001 G2, and C/2001 H1 were detected with both the C2 and C3 coronagraphs, while the remaining comets were visible only with C2. The comets were found at the SOHO website by M. Boschat (C/2001 F2), S. Hoenig (C/2001 G2), T. Scarmato (C/2001 H2), X. Leprette (C/2001 H2, C/2001 H3, C/2001 H4), and M. Oates (remaining comets). C/2001 G2 saturated the C2 images (though the comet was partially obscured by the occulting pylon). C/2001 G3 was very faint.

Comet  2001 UT  α_{2000}  δ_{2000}  MPEC
C/2001 E1 Mar. 15.561  23^{h}44.^{m}7  − 0^{o}11′  2001-F52
C/2001 F2  25.988  0 44.8  + 0 45  2001-F55
C/2001 G2 Apr. 7.540  1 30.5  + 5 10  2001-G34
C/2001 G3  10.729  1 24.6  + 7 20  2001-G34
C/2001 H1  20.096  2 00.5  + 9 31  2001-H32
C/2001 H2  20.479  2 00.4  +10 42  2001-H32
C/2001 H3  20.439  2 00.5  +10 39  2001-H32
C/2001 H4  20.579  2 00.5  +10 41  2001-H32
2001 April 25  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2001 CBAT  Daniel W. E. Green